determine the order and the degree of the given

differential equation; also state whether the equation is linear or nonlinear.

1. t^2\frac{d^2y}{dt^2}+t\frac{dy}{dt}+2y=sin\:t



1
Expert's answer
2021-08-31T11:51:59-0400

"\text{The order of a differential equation is the order of the highest derivative while the }
\\\text{degree is the power of the highest derivative. Given the equation }
 \\t^2\frac{d^2y}{dt^2}+t\frac{dy}{dt}+2y=sin\:t
\\\text{We have that the order is 2 and the degree is 1}"